Circular No. 6790

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997eg IN NGC 5012 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Tsukioka-cho, Toyama, of an apparent supernova (mag 15.6) on ten frames taken around Dec. 4.85 UT and three frames taken on Dec. 5 with a 0.43-m f/6 reflector (+ unfiltered CCD). The limiting magnitude on Dec. 4 was ∼ 18.5. SN 1997eg is located at α = 13^{h}11^{m}36^{s}.73, δ = +22^{o}55′29′′.4 (equinox 2000.0), which is 4′′.1 west and 33′′.1 north of the center of NGC 5012. There is no object at this position on Aoki’s earlier frames taken on fifteen nights between 1996 Apr. 22 and 1997 Aug. 11, or on the Digital Sky Survey. There is a star of mag 13.9 located ∼ 5′′ east of SN 1997eg.

GS 1354–64 W. Heindl, P. Blanco, D. Gruber, M. Pelling, D. MacDonald, D. Mars-

den, and R. Rothschild, Center for Astrophysics and Space Sciences, University of California at San Diego, report: “Pointed observations of GS 1354–64 (IAUC 6772, 6774, 6775, 6779, 6781) with the RXTE/HEXTE on Nov. 22.40, Dec. 5.21, and 12.81 UT yield fluxes of 150, 150, and 130 mCrab (15–200 keV), respectively. In all cases, the spectrum can be described by an exponentially-cutoff power law with photon index 1.0 ± 0.1 and an e-folding energy of ∼ 60 keV. Although GS 1354–64 is in outburst, this spectrum is similar to blackhole candidates in the low (hard) state. The spectrum in the HEXTE band is not consistent with a single power law. However, fitting only in the band 30–100 keV gives photon indices of 2.0 ± 0.1, somewhat harder than the value (2.7) reported with BATSE (IAUC 6774) for these times. Because the source remains active, additional observations at x-ray and other wavelengths continue to be desirable.”

η CARINAE

D. F. Lopes, Observatório Nacional; and A. Damineli, Instituto Astronômico e Geofísico da Universidade de São Paulo, communicate: “η Car has entered a new spectroscopic event phase; CCD spectra collected on Dec. 10 at the coudé focus of the 1.6-m telescope of the Laboratório Nacional de Astrofísica show the complete fading of the high-excitation lines. The excellent agreement between observations and predictions (Damineli et al. 1997, New Astron. 2, 107, and IAUC 6768) supports the hypothesis that the 5.52-yr cycle in these events is periodic.”

1997 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green